OpenStudy (anonymous):

The formula to find the perimeter of a rectangle is P = 2W + 2L. To solve for " W " in this formula, which steps should be followed?

OpenStudy (anonymous):

P -2L = 2W (P - 2L)/2 = W

OpenStudy (anonymous):

2w=P-2l w=(p-2l)/2